# Training Bilingual LMs with Data Constraints in the Targeted Language

### v1

@@ Appendix G Results for Individual datasets for German Target Language Models
No ARC Base DE $21.08\pm 1.19$ $31.52\pm 0.95$ $28.09\pm 0.45$ $56.47\pm 1.16$ $54.80\pm 1.57$ $51.30\pm 1.40$ $40.54$
Base DE $20.73\pm 1.18$ $33.33\pm 0.97$ $28.15\pm 0.45$ $57.40\pm 1.15$ $56.30\pm 1.57$ $50.59\pm 1.41$ $41.09$
Base EN $23.81\pm 1.24$ $41.92\pm 1.01$ $35.73\pm 0.48$ $67.03\pm 1.10$ $66.10\pm 1.50$ $50.75\pm 1.41$ $47.56$
Base EN + OH/ELI5 Filter $25.43\pm 1.27$ $46.17\pm 1.02$ $35.82\pm 0.48$ $67.19\pm 1.10$ $67.50\pm 1.48$ $52.49\pm 1.40$ $49.10$
ARC EN $26.28\pm 1.29$ $46.80\pm 1.02$ $36.44\pm 0.48$ $67.46\pm 1.09$ $70.30\pm 1.45$ $51.38\pm 1.40$ $49.78$
HS EN $23.72\pm 1.24$ $42.42\pm 1.01$ $39.64\pm 0.49$ $70.13\pm 1.07$ $67.20\pm 1.49$ $50.43\pm 1.41$ $48.93$
HS+ARC EN $24.23\pm 1.25$ $44.40\pm 1.02$ $38.51\pm 0.49$ $68.50\pm 1.08$ $68.70\pm 1.47$ $50.43\pm 1.41$ $49.13$
SciQ $26.62\pm 1.29$ $49.33\pm 1.03$ $31.94\pm 0.47$ $63.44\pm 1.12$ $72.40\pm 1.41$ $50.83\pm 1.41$ $49.09$
Inst $23.55\pm 1.24$ $43.27\pm 1.02$ $36.40\pm 0.48$ $67.25\pm 1.09$ $68.90\pm 1.46$ $51.14\pm 1.40$ $48.42$
SciQ+Inst $26.19\pm 1.28$ $46.93\pm 1.02$ $36.02\pm 0.48$ $66.21\pm 1.10$ $73.10\pm 1.40$ $50.67\pm 1.41$ $49.85$
v1 Base EN $21.16\pm 1.19$ $36.83\pm 0.99$ $29.19\pm 0.45$ $60.07\pm 1.14$ $63.50\pm 1.52$ $52.33\pm 1.40$ $43.84$
v2 Base EN $21.16\pm 1.19$ $34.89\pm 0.98$ $29.60\pm 0.46$ $57.45\pm 1.15$ $59.50\pm 1.55$ $50.36\pm 1.41$ $42.16$
v3 Base EN $19.54\pm 1.16$ $35.02\pm 0.98$ $29.46\pm 0.45$ $59.47\pm 1.15$ $61.40\pm 1.54$ $50.51\pm 1.41$ $42.57$
@@
No ARC Base DE $23.48\pm 1.26$ $32.30\pm 0.98$ $30.45\pm 0.48$ $56.09\pm 1.16$ $60.32\pm 1.59$ $52.11\pm 1.45$ $42.46$
Base DE $24.98\pm 1.28$ $34.87\pm 1.00$ $32.28\pm 0.48$ $59.79\pm 1.14$ $61.26\pm 1.58$ $51.52\pm 1.45$ $44.12$
Base EN $23.83\pm 1.26$ $30.66\pm 0.97$ $29.69\pm 0.47$ $56.37\pm 1.16$ $60.84\pm 1.58$ $51.01\pm 1.45$ $42.07$
Base EN + OH/ELI5 Filter $24.27\pm 1.27$ $32.74\pm 0.99$ $29.68\pm 0.47$ $55.88\pm 1.16$ $60.42\pm 1.59$ $51.69\pm 1.45$ $42.45$
ARC EN $23.83\pm 1.26$ $33.05\pm 0.99$ $29.59\pm 0.47$ $56.37\pm 1.16$ $60.11\pm 1.59$ $51.60\pm 1.45$ $42.43$
HS EN $23.75\pm 1.26$ $31.81\pm 0.98$ $30.32\pm 0.47$ $57.40\pm 1.15$ $59.68\pm 1.59$ $51.52\pm 1.45$ $42.41$
HS+ARC EN $24.54\pm 1.28$ $33.01\pm 0.99$ $30.10\pm 0.47$ $55.01\pm 1.16$ $60.11\pm 1.59$ $50.84\pm 1.45$ $42.27$
SciQ $23.66\pm 1.26$ $32.21\pm 0.98$ $28.84\pm 0.47$ $56.26\pm 1.16$ $61.58\pm 1.58$ $52.96\pm 1.45$ $42.58$
Inst $25.42\pm 1.29$ $32.79\pm 0.99$ $29.59\pm 0.47$ $55.93\pm 1.16$ $61.05\pm 1.58$ $50.68\pm 1.45$ $42.58$
SciQ+Inst $23.22\pm 1.25$ $33.67\pm 0.99$ $29.42\pm 0.47$ $55.93\pm 1.16$ $61.68\pm 1.58$ $51.27\pm 1.45$ $42.53$
v1 Base EN $23.92\pm 1.27$ $35.27\pm 1.01$ $32.79\pm 0.49$ $59.74\pm 1.14$ $62.42\pm 1.57$ $50.68\pm 1.45$ $44.14$
v2 Base EN $24.27\pm 1.27$ $36.19\pm 1.01$ $32.72\pm 0.48$ $59.41\pm 1.15$ $63.89\pm 1.56$ $52.53\pm 1.45$ $44.84$
v3 Base EN $23.83\pm 1.26$ $36.19\pm 1.01$ $34.06\pm 0.49$ $61.04\pm 1.14$ $63.26\pm 1.56$ $51.77\pm 1.45$ $45.03$
RPJv2 $24.10\pm 1.27$ $33.10\pm 0.99$ $30.06\pm 0.47$ $55.44\pm 1.16$ $59.05\pm 1.60$ $50.08\pm 1.45$ $41.97$
RefinedWeb $25.33\pm 1.29$ $32.12\pm 0.98$ $30.56\pm 0.48$ $57.34\pm 1.15$ $62.32\pm 1.57$ $51.77\pm 1.45$ $43.24$
FineWeb $24.45\pm 1.28$ $33.67\pm 0.99$ $31.20\pm 0.48$ $56.31\pm 1.16$ $62.32\pm 1.57$ $51.01\pm 1.45$ $43.16$
FineWebEDU $25.59\pm 1.29$ $35.00\pm 1.00$ $31.19\pm 0.48$ $56.86\pm 1.16$ $62.32\pm 1.57$ $51.52\pm 1.45$ $43.75$
Table 13: Evaluation of $300$M parameter XL model on “General Understanding Tasks” focusing on general reasoning, language understanding, and science knowledge in translated German. Results show the length normalized accuracy for individual datasets and the average over all datasets for all datasets.
@@
No ARC Base DE $21.33\pm 1.20$ $35.40\pm 0.98$ $30.88\pm 0.46$ $59.14\pm 1.15$ $60.50\pm 1.55$ $50.67\pm 1.41$ $42.99$
Base DE $22.78\pm 1.23$ $37.92\pm 1.00$ $33.06\pm 0.47$ $62.35\pm 1.13$ $64.70\pm 1.51$ $51.14\pm 1.40$ $45.33$
Base EN $25.94\pm 1.28$ $48.15\pm 1.03$ $48.50\pm 0.50$ $71.55\pm 1.05$ $73.10\pm 1.40$ $52.01\pm 1.40$ $53.21$
Base EN + OH/ELI5 Filter $29.52\pm 1.33$ $57.03\pm 1.02$ $49.96\pm 0.50$ $71.87\pm 1.05$ $77.90\pm 1.31$ $53.83\pm 1.40$ $56.69$
ARC EN $31.83\pm 1.36$ $57.24\pm 1.02$ $48.80\pm 0.50$ $73.07\pm 1.04$ $77.00\pm 1.33$ $52.72\pm 1.40$ $56.78$
HS EN $28.84\pm 1.32$ $49.83\pm 1.03$ $54.83\pm 0.50$ $75.14\pm 1.01$ $75.10\pm 1.37$ $54.85\pm 1.40$ $56.43$
HS+ARC EN $28.84\pm 1.32$ $56.10\pm 1.02$ $53.04\pm 0.50$ $73.67\pm 1.03$ $77.90\pm 1.31$ $54.38\pm 1.40$ $57.32$
SciQ $30.63\pm 1.35$ $57.28\pm 1.02$ $39.79\pm 0.49$ $68.28\pm 1.09$ $80.40\pm 1.26$ $53.91\pm 1.40$ $55.05$
Inst $27.22\pm 1.30$ $53.66\pm 1.02$ $50.30\pm 0.50$ $71.49\pm 1.05$ $76.20\pm 1.35$ $53.51\pm 1.40$ $55.40$
SciQ+Inst $29.10\pm 1.33$ $55.22\pm 1.02$ $48.34\pm 0.50$ $71.71\pm 1.05$ $78.50\pm 1.30$ $53.04\pm 1.40$ $55.98$
v1 Base EN $20.73\pm 1.18$ $38.97\pm 1.00$ $36.09\pm 0.48$ $63.33\pm 1.12$ $65.40\pm 1.51$ $52.01\pm 1.40$ $46.09$
v2 Base EN $22.70\pm 1.22$ $38.97\pm 1.00$ $36.55\pm 0.48$ $64.80\pm 1.11$ $69.30\pm 1.46$ $51.22\pm 1.40$ $47.26$
v3 Base EN $20.65\pm 1.18$ $40.36\pm 1.01$ $36.57\pm 0.48$ $63.66\pm 1.12$ $70.50\pm 1.44$ $51.54\pm 1.40$ $47.21$
RPJv2 $26.96\pm 1.30$ $50.42\pm 1.03$ $51.10\pm 0.50$ $70.57\pm 1.06$ $77.60\pm 1.32$ $55.64\pm 1.40$ $55.38$
RefinedWeb $27.90\pm 1.31$ $54.59\pm 1.02$ $54.91\pm 0.50$ $73.23\pm 1.03$ $77.60\pm 1.32$ $56.35\pm 1.39$ $57.43$
FineWeb $27.82\pm 1.31$ $52.15\pm 1.03$ $56.24\pm 0.50$ $73.94\pm 1.02$ $74.40\pm 1.38$ $55.72\pm 1.40$ $56.71$
FineWebEDU $38.14\pm 1.42$ $66.37\pm 0.97$ $54.88\pm 0.50$ $72.25\pm 1.04$ $84.60\pm 1.14$ $56.04\pm 1.39$ $62.05$
Table 14: Evaluation of $1$B parameter XL model on “General Understanding Tasks” focusing on general reasoning, language understanding, and science knowledge in English. Results show the length normalized accuracy for individual datasets and the average over all datasets for all datasets.
@@
No ARC Base DE $25.07\pm 1.29$ $36.95\pm 1.02$ $35.64\pm 0.49$ $59.74\pm 1.14$ $64.53\pm 1.55$ $52.62\pm 1.45$ $45.76$
Base DE $27.44\pm 1.32$ $38.81\pm 1.03$ $39.53\pm 0.51$ $63.17\pm 1.13$ $67.68\pm 1.52$ $52.62\pm 1.45$ $48.21$
Base EN $25.07\pm 1.29$ $37.70\pm 1.02$ $36.18\pm 0.50$ $59.36\pm 1.15$ $63.05\pm 1.57$ $50.51\pm 1.45$ $45.31$
Base EN + OH/ELI5 Filter $25.77\pm 1.30$ $40.00\pm 1.03$ $35.94\pm 0.50$ $59.25\pm 1.15$ $65.05\pm 1.55$ $51.18\pm 1.45$ $46.20$
ARC EN $26.91\pm 1.32$ $39.34\pm 1.03$ $35.77\pm 0.50$ $58.65\pm 1.15$ $67.58\pm 1.52$ $52.11\pm 1.45$ $46.73$
HS EN $24.89\pm 1.28$ $36.90\pm 1.02$ $37.11\pm 0.50$ $60.77\pm 1.14$ $63.47\pm 1.56$ $52.87\pm 1.45$ $46.00$
HS+ARC EN $26.74\pm 1.31$ $38.50\pm 1.02$ $36.71\pm 0.50$ $59.96\pm 1.14$ $63.89\pm 1.56$ $51.35\pm 1.45$ $46.19$
SciQ $27.35\pm 1.32$ $39.38\pm 1.03$ $33.23\pm 0.49$ $58.76\pm 1.15$ $64.21\pm 1.56$ $51.18\pm 1.45$ $45.69$
Inst $25.59\pm 1.29$ $38.14\pm 1.02$ $36.22\pm 0.50$ $60.01\pm 1.14$ $64.74\pm 1.55$ $51.01\pm 1.45$ $45.95$
SciQ+Inst $25.24\pm 1.29$ $39.38\pm 1.03$ $35.34\pm 0.49$ $59.74\pm 1.14$ $64.21\pm 1.56$ $52.70\pm 1.45$ $46.10$
v1 Base EN $26.12\pm 1.30$ $40.93\pm 1.03$ $40.39\pm 0.51$ $61.48\pm 1.14$ $67.58\pm 1.52$ $51.94\pm 1.45$ $48.07$
v2 Base EN $25.51\pm 1.29$ $42.30\pm 1.04$ $40.75\pm 0.51$ $62.19\pm 1.13$ $71.37\pm 1.47$ $52.45\pm 1.45$ $49.09$
v3 Base EN $26.21\pm 1.30$ $40.84\pm 1.03$ $43.08\pm 0.51$ $64.09\pm 1.12$ $66.74\pm 1.53$ $51.52\pm 1.45$ $48.75$
RPJv2 $24.71\pm 1.28$ $37.17\pm 1.02$ $36.53\pm 0.50$ $58.65\pm 1.15$ $65.26\pm 1.55$ $52.70\pm 1.45$ $45.84$
RefinedWeb $25.42\pm 1.29$ $38.81\pm 1.03$ $38.25\pm 0.50$ $58.38\pm 1.15$ $64.21\pm 1.56$ $51.01\pm 1.45$ $46.01$
FineWeb $26.12\pm 1.30$ $36.90\pm 1.02$ $38.09\pm 0.50$ $59.85\pm 1.14$ $64.11\pm 1.56$ $53.89\pm 1.45$ $46.49$
FineWebEDU $26.91\pm 1.32$ $42.39\pm 1.04$ $37.40\pm 0.50$ $60.45\pm 1.14$ $65.37\pm 1.54$ $50.42\pm 1.45$ $47.16$
Table 15: Evaluation of $1$B parameter XL model on “General Understanding Tasks” focusing on general reasoning, language understanding, and science knowledge in translated German. Results show the length normalized accuracy for individual datasets and the average over all datasets for all datasets.

### v2

@@ Appendix I Results for Individual datasets for German Target Language Models
Small DE $21.33\pm 1.20$ $27.90\pm 0.92$ $26.20\pm 0.44$ $51.20\pm 1.17$ $42.70\pm 1.56$ $49.72\pm 1.41$ $36.51$
No ARC Large DE $21.08\pm 1.19$ $31.52\pm 0.95$ $28.09\pm 0.45$ $56.47\pm 1.16$ $54.80\pm 1.57$ $51.30\pm 1.40$ $40.54$
Large DE $20.73\pm 1.18$ $33.33\pm 0.97$ $28.15\pm 0.45$ $57.40\pm 1.15$ $56.30\pm 1.57$ $50.59\pm 1.41$ $41.09$
Small DE + Large EN $23.81\pm 1.24$ $41.92\pm 1.01$ $35.73\pm 0.48$ $67.03\pm 1.10$ $66.10\pm 1.50$ $50.75\pm 1.41$ $47.56$
Small DE + Large EN DCLM Filter $25.43\pm 1.27$ $46.17\pm 1.02$ $35.82\pm 0.48$ $67.19\pm 1.10$ $67.50\pm 1.48$ $52.49\pm 1.40$ $49.10$
Small DE + Large HS EN $23.72\pm 1.24$ $42.42\pm 1.01$ $39.64\pm 0.49$ $70.13\pm 1.07$ $67.20\pm 1.49$ $50.43\pm 1.41$ $48.93$
Small DE +Large ARC EN $26.28\pm 1.29$ $46.80\pm 1.02$ $36.44\pm 0.48$ $67.46\pm 1.09$ $70.30\pm 1.45$ $51.38\pm 1.40$ $49.78$
+ Large S EN $26.62\pm 1.29$ $49.33\pm 1.03$ $31.94\pm 0.47$ $63.44\pm 1.12$ $72.40\pm 1.41$ $50.83\pm 1.41$ $49.09$
Small DE + Large S+I EN $26.19\pm 1.28$ $46.93\pm 1.02$ $36.02\pm 0.48$ $66.21\pm 1.10$ $73.10\pm 1.40$ $50.67\pm 1.41$ $49.85$
Small DE + Large v1 $21.16\pm 1.19$ $36.83\pm 0.99$ $29.19\pm 0.45$ $60.07\pm 1.14$ $63.50\pm 1.52$ $52.33\pm 1.40$ $43.84$
Small DE + Large v2 $21.16\pm 1.19$ $34.89\pm 0.98$ $29.60\pm 0.46$ $57.45\pm 1.15$ $59.50\pm 1.55$ $50.36\pm 1.41$ $42.16$
Small DE + Large v3 $19.54\pm 1.16$ $35.02\pm 0.98$ $29.46\pm 0.45$ $59.47\pm 1.15$ $61.40\pm 1.54$ $50.51\pm 1.41$ $42.57$
@@
Small DE $23.83\pm 1.26$ $30.71\pm 0.97$ $28.33\pm 0.47$ $53.92\pm 1.16$ $47.58\pm 1.62$ $52.20\pm 1.45$ $39.43$
No ARC Large DE $23.48\pm 1.26$ $32.30\pm 0.98$ $30.45\pm 0.48$ $56.09\pm 1.16$ $60.32\pm 1.59$ $52.11\pm 1.45$ $42.46$
Large DE $24.98\pm 1.28$ $34.87\pm 1.00$ $32.28\pm 0.48$ $59.79\pm 1.14$ $61.26\pm 1.58$ $51.52\pm 1.45$ $44.12$
Small DE + Large EN $23.83\pm 1.26$ $30.66\pm 0.97$ $29.69\pm 0.47$ $56.37\pm 1.16$ $60.84\pm 1.58$ $51.01\pm 1.45$ $42.07$
Small DE + Large EN DCLM Filter $24.27\pm 1.27$ $32.74\pm 0.99$ $29.68\pm 0.47$ $55.88\pm 1.16$ $60.42\pm 1.59$ $51.69\pm 1.45$ $42.45$
Small DE + Large HS EN $23.75\pm 1.26$ $31.81\pm 0.98$ $30.32\pm 0.47$ $57.40\pm 1.15$ $59.68\pm 1.59$ $51.52\pm 1.45$ $42.41$
Small DE +Large ARC EN $23.83\pm 1.26$ $33.05\pm 0.99$ $29.59\pm 0.47$ $56.37\pm 1.16$ $60.11\pm 1.59$ $51.60\pm 1.45$ $42.43$
Small DE + Large S EN $23.66\pm 1.26$ $32.21\pm 0.98$ $28.84\pm 0.47$ $56.26\pm 1.16$ $61.58\pm 1.58$ $52.96\pm 1.45$ $42.58$
Small DE + Large S+I EN $23.22\pm 1.25$ $33.67\pm 0.99$ $29.42\pm 0.47$ $55.93\pm 1.16$ $61.68\pm 1.58$ $51.27\pm 1.45$ $42.53$
Small DE + Large v1 $23.92\pm 1.27$ $35.27\pm 1.01$ $32.79\pm 0.49$ $59.74\pm 1.14$ $62.42\pm 1.57$ $50.68\pm 1.45$ $44.14$
Small DE + Large v2 $24.27\pm 1.27$ $36.19\pm 1.01$ $32.72\pm 0.48$ $59.41\pm 1.15$ $63.89\pm 1.56$ $52.53\pm 1.45$ $44.84$
Small DE + Large v3 $23.83\pm 1.26$ $36.19\pm 1.01$ $34.06\pm 0.49$ $61.04\pm 1.14$ $63.26\pm 1.56$ $51.77\pm 1.45$ $45.03$
Small DE + Large RPJv2 $24.10\pm 1.27$ $33.10\pm 0.99$ $30.06\pm 0.47$ $55.44\pm 1.16$ $59.05\pm 1.60$ $50.08\pm 1.45$ $41.97$
Small DE + Large RFW $25.33\pm 1.29$ $32.12\pm 0.98$ $30.56\pm 0.48$ $57.34\pm 1.15$ $62.32\pm 1.57$ $51.77\pm 1.45$ $43.24$
Small DE + Large FWE $25.59\pm 1.29$ $35.00\pm 1.00$ $31.19\pm 0.48$ $56.86\pm 1.16$ $62.32\pm 1.57$ $51.52\pm 1.45$ $43.75$
Table 13: Evaluation of $300$M parameter model on “General Understanding Tasks” focusing on general reasoning, language understanding, and science knowledge in translated German. Results show the length normalized accuracy for individual datasets and the average over all datasets for all datasets.
@@
Small DE $20.73\pm 1.18$ $27.53\pm 0.92$ $25.94\pm 0.44$ $52.88\pm 1.16$ $36.80\pm 1.53$ $47.12\pm 1.40$ $35.17$
No ARC Large DE $21.33\pm 1.20$ $35.40\pm 0.98$ $30.88\pm 0.46$ $59.14\pm 1.15$ $60.50\pm 1.55$ $50.67\pm 1.41$ $42.99$
Large DE $22.78\pm 1.23$ $37.92\pm 1.00$ $33.06\pm 0.47$ $62.35\pm 1.13$ $64.70\pm 1.51$ $51.14\pm 1.40$ $45.33$
Small DE + Large EN $25.94\pm 1.28$ $48.15\pm 1.03$ $48.50\pm 0.50$ $71.55\pm 1.05$ $73.10\pm 1.40$ $52.01\pm 1.40$ $53.21$
Small DE + Large EN DCLM Filter $29.52\pm 1.33$ $57.03\pm 1.02$ $49.96\pm 0.50$ $71.87\pm 1.05$ $77.90\pm 1.31$ $53.83\pm 1.40$ $56.69$
Small DE + Large HS EN $28.84\pm 1.32$ $49.83\pm 1.03$ $54.83\pm 0.50$ $75.14\pm 1.01$ $75.10\pm 1.37$ $54.85\pm 1.40$ $56.43$
Small DE +Large ARC EN $31.83\pm 1.36$ $57.24\pm 1.02$ $48.80\pm 0.50$ $73.07\pm 1.04$ $77.00\pm 1.33$ $52.72\pm 1.40$ $56.78$
Small DE + Large S EN $30.63\pm 1.35$ $57.28\pm 1.02$ $39.79\pm 0.49$ $68.28\pm 1.09$ $80.40\pm 1.26$ $53.91\pm 1.40$ $55.05$
Small DE + Large S+I EN $29.10\pm 1.33$ $55.22\pm 1.02$ $48.34\pm 0.50$ $71.71\pm 1.05$ $78.50\pm 1.30$ $53.04\pm 1.40$ $55.98$
Small DE + Large v1 $20.73\pm 1.18$ $38.97\pm 1.00$ $36.09\pm 0.48$ $63.33\pm 1.12$ $65.40\pm 1.51$ $52.01\pm 1.40$ $46.09$
Small DE + Large v2 $22.70\pm 1.22$ $38.97\pm 1.00$ $36.55\pm 0.48$ $64.80\pm 1.11$ $69.30\pm 1.46$ $51.22\pm 1.40$ $47.26$
Small DE + Large v3 $20.65\pm 1.18$ $40.36\pm 1.01$ $36.57\pm 0.48$ $63.66\pm 1.12$ $70.50\pm 1.44$ $51.54\pm 1.40$ $47.21$
Small DE + Large RPJv2 $26.96\pm 1.30$ $50.42\pm 1.03$ $51.10\pm 0.50$ $70.57\pm 1.06$ $77.60\pm 1.32$ $55.64\pm 1.40$ $55.38$
Small DE + Large RFW $27.90\pm 1.31$ $54.59\pm 1.02$ $54.91\pm 0.50$ $73.23\pm 1.03$ $77.60\pm 1.32$ $56.35\pm 1.39$ $57.43$
Small DE + Large FWE $38.14\pm 1.42$ $66.37\pm 0.97$ $54.88\pm 0.50$ $72.25\pm 1.04$ $84.60\pm 1.14$ $56.04\pm 1.39$ $62.05$
Table 14: Evaluation of $1.3$B parameter model on “General Understanding Tasks” focusing on general reasoning, language understanding, and science knowledge in English. Results show the length normalized accuracy for individual datasets and the average over all datasets for all datasets.
@@
Small DE $22.69\pm 1.24$ $30.31\pm 0.97$ $28.99\pm 0.47$ $54.84\pm 1.16$ $40.00\pm 1.59$ $50.42\pm 1.45$ $37.88$
No ARC Large DE $25.07\pm 1.29$ $36.95\pm 1.02$ $35.64\pm 0.49$ $59.74\pm 1.14$ $64.53\pm 1.55$ $52.62\pm 1.45$ $45.76$
Large DE $27.44\pm 1.32$ $38.81\pm 1.03$ $39.53\pm 0.51$ $63.17\pm 1.13$ $67.68\pm 1.52$ $52.62\pm 1.45$ $48.21$
Small DE + Large EN $25.07\pm 1.29$ $37.70\pm 1.02$ $36.18\pm 0.50$ $59.36\pm 1.15$ $63.05\pm 1.57$ $50.51\pm 1.45$ $45.31$
Small DE + Large EN DCLM Filter $25.77\pm 1.30$ $40.00\pm 1.03$ $35.94\pm 0.50$ $59.25\pm 1.15$ $65.05\pm 1.55$ $51.18\pm 1.45$ $46.20$
Small DE + Large HS EN $24.89\pm 1.28$ $36.90\pm 1.02$ $37.11\pm 0.50$ $60.77\pm 1.14$ $63.47\pm 1.56$ $52.87\pm 1.45$ $46.00$
Small DE +Large ARC EN $26.91\pm 1.32$ $39.34\pm 1.03$ $35.77\pm 0.50$ $58.65\pm 1.15$ $67.58\pm 1.52$ $52.11\pm 1.45$ $46.73$
+ Large S EN $27.35\pm 1.32$ $39.38\pm 1.03$ $33.23\pm 0.49$ $58.76\pm 1.15$ $64.21\pm 1.56$ $51.18\pm 1.45$ $45.69$
Small DE + Large S+I EN $25.24\pm 1.29$ $39.38\pm 1.03$ $35.34\pm 0.49$ $59.74\pm 1.14$ $64.21\pm 1.56$ $52.70\pm 1.45$ $46.10$
Small DE + Large v1 $26.12\pm 1.30$ $40.93\pm 1.03$ $40.39\pm 0.51$ $61.48\pm 1.14$ $67.58\pm 1.52$ $51.94\pm 1.45$ $48.07$
Small DE + Large v2 $25.51\pm 1.29$ $42.30\pm 1.04$ $40.75\pm 0.51$ $62.19\pm 1.13$ $71.37\pm 1.47$ $52.45\pm 1.45$ $49.09$
Small DE + Large v3 $26.21\pm 1.30$ $40.84\pm 1.03$ $43.08\pm 0.51$ $64.09\pm 1.12$ $66.74\pm 1.53$ $51.52\pm 1.45$ $48.75$
Small DE + Large RPJv2 $24.71\pm 1.28$ $37.17\pm 1.02$ $36.53\pm 0.50$ $58.65\pm 1.15$ $65.26\pm 1.55$ $52.70\pm 1.45$ $45.84$
Small DE + Large RFW $25.42\pm 1.29$ $38.81\pm 1.03$ $38.25\pm 0.50$ $58.38\pm 1.15$ $64.21\pm 1.56$ $51.01\pm 1.45$ $46.01$
Small DE + Large FWE $26.91\pm 1.32$ $42.39\pm 1.04$ $37.40\pm 0.50$ $60.45\pm 1.14$ $65.37\pm 1.54$ $50.42\pm 1.45$ $47.16$
Table 15: Evaluation of $1.3$B parameter model on “General Understanding Tasks” focusing on general reasoning, language understanding, and science knowledge in translated German. Results show the length normalized accuracy for individual datasets and the average over all datasets for all datasets.